How would you write and exponential function whose graph passes through (0,0.4) and (2,6.4)?
Answer by stanbon(75887)

How would you write and exponential function whose graph passes through (0,0.4) and (2,6.4)?
----------------
Form y=ab^x
0.4 = ab^0
a=0.4
-------
y=0.4b^x
6.4=0.4b^2
16=b^2
b=4
------------------
EQUATION:
y=0.4(4)^x is the exponential equation passing thru the two points.
====================
Cheers,
Stan H.